# Enhancing Repository-Level Code Generation with Integrated Contextual Information

### v1

@@ 5.2. RQ2: Ablation Study
Table 4. Evaluation results of different variants of our approach on two benchmarks, using CodeLlama-13B-Instruct
Java Rust
pass@k (%) pass@k (%)
Variant $k=1$ $k=3$ $k=5$ $k=1$ $k=3$ $k=5$
FULL 44.67 53.97 57.55 52.67 62.00 65.89
-CR 27.54 38.84 43.66 23.33 35.40 41.04
-TC 39.50 49.48 53.38 49.89 59.61 63.80
@@
We evaluate the above three variants on the same benchmarks as RQ1, also using CodeLlama-13B-Instruct as the underlying LLM. We compare the incomplete ones with FULL, and the results are illustrated in Table 4. We can conclude that CatCoder benefits from the two components, as the pass@$k$ scores of both incomplete variants drop noticeably. In detail, for the -CR variant, the pass@$k$ scores decrease by up to 38.35% on the Java benchmark and by up to 56.03% on the Rust benchmark. And for the -TC variant, the pass@$k$ scores drop by up to 11.57% on the Java benchmark and by up to 5.3% on the Rust benchmark.
@@ 5.3. RQ3: Generalizability
Figure 6. Evaluation results of the In-File and CatCoder methods on the Java benchmark, using various large language models. The model acronyms on the $x$-axis corresponding to the ones in Table 2. The results are sorted in ascending order based on the values of CatCoder, and the “Average” bar shows the average pass@1 score of all evaluated models.
@@ 5.3.1. Results on Generalizability
As described in Section 4.4.2, we select a variety of LLMs for the generalizablity study, and the details of the models are listed in Table 2. Figure 6 shows the pass@$1$ scores of CatCoder on the Java benchmark using the selected LLMs. From the plot, we can notice that CatCoder is able to improve the performance of every model, compared with the In-File context. On average, pass@$1$ scores increase by 32.9%. Therefore, we can conclude that CatCoder can be generalized to different LLMs.
@@ 5.3.2. Comparison of different LLMs
We can also learn some facts from pass@$1$ scores across different LLMs [^10] Here, we use the model acronyms in Table 2 to refer to the models.:
@@
❶ Although the selected general purpose LLMs claim to be powerful in many aspects (lla, 2024), they fall behind all code-specialized models except for CG-7B in our benchmarks. One example is ML-8B versus CL-7B. Although CL-7B is fine-tuned from an older foundation model, i.e., Llama 2 (Touvron et al., 2023), it outperforms the new model ML-8B (Llama 3 (lla, 2024)). This indicates that code generation tasks benefit a lot from fine-tuning models using high-quality code datasets.
@@
❷ From the trend in Figure 6, although larger models tend to have a better absolute performance than smaller models, the improvement by CatCoder is more on smaller models. For example, the pass@$1$ scores of CG-7B, ML-8B, CQ-7B, DS-6.7B increase more, compared with SC-15B and CL-34B. This fact provides an insight that we can deploy smaller LLMs on personal devices and use CatCoder to keep up with larger models.
❸ In addition, we can notice that some smaller models even achieve higher performance than larger models. For example, as shown in Figure 6, the performance of two smaller models (i.e., CQ-7B and DS-6.7B) outperforms SC-15B while SC-15B has over twice amount of parameters. This highlights the importance of model architecture and training datasets.
@@
Answer to RQ3: CatCoder is generalizable to many different LLMs, as it improves the pass@$1$ scores of all selected LLMs in our evaluation. In addition, the performance of different LLMs on the benchmark provides valuable insights.
@@ 6. Discussion
In this section, we discuss the limitation of our approach CatCoder and the threats to validity.
6.1. When CatCoder fails
@@
Figure 7 shows an example of a failure case generated by CatCoder. The generated code uses all APIs correctly and looks similar to the ground truth code but fails to pass some unit tests because it lacks the logic to handle the null pointers (i.e., the green part in Figure 7). From this case, we can see that function code generation is a complex task, and successful usage of repository context does not always lead to correct code. CatCoder is limited in improving the logical reasoning ability of LLMs, and possible future work could be fine-tuning LLMs with code repositories or using reinforcement learning-based approaches (Le et al., 2022).
@@ 6.2. Threats to Validity
6.2.1. Support for More Languages
@@
6.2.2. Potential Data Leakage
Our experiments are conducted on the latest state-of-the-art LLMs, and their knowledge cutoff dates are very recent. It is possible that the code repositories in our benchmarks have been included in the training corpora. Thus, the potential data leakage problem could affect the rigor of our results and findings. For example, in this case, it could make the pass@$k$ scores of all approaches higher, as the LLM could output memorized code without having to understand the repository context. In the future, possible improvements may include:
  1. (1)
    Replacing the benchmarks with the most up-to-date ones (Liu et al., 2023; Li et al., 2024) and re-evaluating the approach.
  2. (2)
    Refactoring the benchmarks by renaming the identifiers in the dataset repositories without changing the semantics (Zan et al., 2022).
6.2.3. Efficiency Issue
As CatCoder requires the retrieval and analysis of the source code, there is a noticeable latency from input to output, especially when processing large repositories. Thus, CatCoder is less efficient than commercial tools such as Github Copilot. However, compared to methods that require multiple iterations of model generation, CatCoder’s efficiency is still competitive. In the future, we plan to mitigate this issue by making the components concurrent and caching intermediate data (e.g., embedding vectors and results of type analysis).
@@ 7.1. Large Language Models for Code
Large language models for code (i.e. code LLMs) refer to a family of Transformer (Vaswani et al., 2017) models pre-trained or fine-tuned on massive code datasets. Previous work has shown the exceptional performance of code LLMs in various code-related tasks, such as code generation (Chen et al., 2021), tese case generation (Tufano et al., 2020) and program repair (Xia et al., 2023). In general, pre-trained language models for code can be categorized into the following three types:
@@
    Encoder-only models, such as CodeBERT (Feng et al., 2020) and GraphCodeBERT (Guo et al., 2020). These models are pre-trained using Masked Language Modeling (MLM) techniques, and are pupular and effective for tasks like classification, where a deep understanding of input data is required.
@@
    Encoder-Decoder models, such as the CodeT5 family (Wang et al., 2021, 2023). These models first encode the input into an internal representation and then decode this representation into an output sequence.
@@
    Decoder-only models, such as Codex (Chen et al., 2021), CodeGen (Nijkamp et al., 2022), StarCoder (Li et al., 2023) and CodeLlama (Roziere et al., 2023). These models are pre-trained using Next Token Prediction tasks, and are powerful in generative tasks.
@@ 7.2. Benchmarks for Repository-Level Code Generation
Iyer et al. (Iyer et al., 2018) are among the first to highlight the importance of the programmatic context and introduce the task of generating class member functions with natural language descriptions and class contexts. In summary, they construct Concode, a large training dataset consisting of Java classes from online code repositories, and propose an encoder-decoder model for code generation. With the development of large language models, many subsequent repository-level code generation benchmarks arise in academia. Du et al. (Du et al., 2023) make an attempt to evaluate LLMs in the more challenging class-level code generation scenario and manually construct a benchmark named ClassEval, which consists of 100 class-level Python code generation tasks.
Compared with previous benchmarks, for example HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), ClassEval is more pragmatic. However, it is still limited because, although the methods in the tasks are not standalone, the programmatic contexts are still restricted to very common libraries and methods within the same file. In real-world scenarios, a code repository generally spans a large number of files, making cross-file contexts essential for LLMs to generate the correct code (Ding et al., 2024).
Toward more realistic evaluation, subsequent benchmarks such as CoderEval (Yu et al., 2024), CrossCodeEval (Ding et al., 2024) and CoderUJB (Zeng et al., 2024) are proposed, all of which are built on a set of real-world open-source repositories and utilize cross-file context for evaluation. A key novelty of CoderEval and CrossCodeEval lies in their ability to support multilingual code generation tasks, while CoderUJB innovatively spans five kinds of practical programming tasks.
As LLMs evolve, Liu et al. (Liu et al., 2023) point out data leakage and memorization problems, which may affect the integrity and trustworthiness of model evaluation results. To mitigate this issue, they propose RepoBench and regularly update the dataset to keep code samples up-to-date. Similarly, Li et al. (Li et al., 2024) propose EvoCodeBench, which includes an automatic pipeline to evolve itself.
@@ 7.3. Repository-Level Code Generation with LLMs
Many existing approaches for repository-level code generation perform retrieval. Some work, such as (Zan et al., 2022; Lu et al., 2022), searches for useful information from documentation or external databases. Specifically, Zan et al. (Zan et al., 2022) train a dual-encoder model, which encodes the natural language descriptions and API information, respectively, to retrieve possible APIs from the documentation of the library. Lu et al. (Lu et al., 2022) propose a retrieval-augmented generation approach named ReAcc, which retrieves similar code from a code database they built. However, these approaches are not flexible and general as their retrieval sources (e.g., API documentation and external databases) may not be available in other repositories.
Other retrieval-based approaches explore retrieving similar code snippets within the repositories. Shrivastava et al. (Shrivastava et al., 2023) propose a framework named Repo-Level Prompt Generator to complete the code for holes in single lines. In the framework, they train a prompt proposal classifier that takes the repository files and a set of prompt proposals and outputs the predicted prompt proposal. Zhang et al. (Zhang et al., 2023) propose RepoCoder, which utilizes repository information by iteratively retrieving similar code. In the initial iteration, it uses the natural language descriptions as query input, and in subsequent iterations, it uses previously generated code for query.
Our work CatCoder combines retrieval-based techniques with a new kind of information called type context, and the comparison with RepoCoder in Section 5 shows that type context serves as essential auxillary information for statically type languages. In addition, CatCoder does not require model training or multiple generation iterations, making it more practical in real-world settings.
In addition to the above work, which mainly performs prompt engineering, several model training frameworks (Wu et al., 2024; Guo et al., 2024) and decoding strategies (Agrawal et al., 2024) have been proposed to support repository-level code generation. As our approach CatCoder does not modify the underlying LLM, it can be combined with these frameworks to improve effectiveness.

### v2

@@ 5.2. RQ2: Ablation Study
Table 7. Evaluation results of different variants on two benchmarks, using CodeLlama-13B-Instruct
Variant Java Rust
compile@k (%) pass@k (%) compile@k (%) pass@k (%)
11 33 55 11 33 55 11 33 55 11 33 55
FULL 71.9 84.8 88.3 44.7 54.0 57.5 65.1 75.5 78.9 52.7 62.0 65.9
-CR 57.9 75.4 81.4 27.5 38.8 43.7 35.0 50.7 56.3 23.3 35.4 41.0
-TC 67.1 80.3 84.8 39.5 49.5 53.4 59.6 70.1 73.8 49.9 59.6 63.8
@@
We evaluate the above three variants on the same benchmarks as RQ1, also using CodeLlama-13B-Instruct as the underlying LLM. We compare the incomplete ones with FULL, and the results are illustrated in Table 7. We can conclude that CatCoder benefits from the two components, as the compile@$k$ and pass@$k$ scores of both incomplete variants drop noticeably. In detail, for the -CR variant, the compile@$k$ and pass@$k$ scores decrease by up to 19.43% and 38.35%, respectively, on the Java benchmark, and by up to 46.24% and 56.03% on the Rust benchmark. And for the -TC variant, the compile@$k$ and pass@$k$ scores drop by up to 6.63% and 11.57% respectively on the Java benchmark, and by up to 8.52% and 5.3% on the Rust benchmark.
Notably, removing Type Context Extraction leads to a smaller performance decline compared to removing Relevant Code Retrieval. This observation underscores the effectiveness of our hybrid retrieval mechanism. However, we emphasize that the importance of type context is not diminished. In many instances, particularly those involving functionalities that are symmetric to, or minor variations of, existing code in the repository, the retriever is able to find highly relevant examples. The presence of duplicated code snippets benefits retrieval-based generation by providing the LLM with strong, nearly complete templates for the target function. In such cases, the retrieved code is sufficient for the LLM to produce a correct solution, rendering the additional type context less critical. However, the performance gap between the -TC variant and FULL highlights a subset of more challenging problems where retrieval alone is insufficient. The qualitative examples in Fig. 5 illustrate a typical failure mode: the retrieved code may include distracting snippets or reference elements outside the current scope, resulting in generation errors. Therefore, we argue that our Type Context Extraction component is an essential and complementary mechanism.
@@ 5.3. RQ3: Generalizability
Figure 6. pass@1 performance comparison using various LLMs on the Java and Rust benchmarks. Subfigures (a) and (b) compare CatCoder with baseline methods—In-File and RepoCoder—while (c) and (d) compare CatCoder with its variants. Model acronyms on the $x$-axis are consistent with those listed in Table 4, and variant names refer to those introduced in RQ2. Models are sorted in ascending order based on the performance of CatCoder. The “Average” bar represents the mean pass@1 score across all evaluated models.
To assess whether the benefits of CatCoder are specific to a single LLM or generalize across different models, we conduct an extensive generalizability study. As described in Section 4.4.2, we select a diverse set of LLMs, including both code-specialized and general-purpose models of varying sizes (Table 4). We then evaluate the performance of the In-File, RepoCoder, and CatCoder approaches, as well as the CatCoder variants (-CR and -TC), using each of these models on both our Java and Rust benchmarks.
@@ 5.3.1. Results on Generalizability
Fig. 6 shows the pass@$1$ scores for these experiments. The results for both Java (Fig. 6a, 6c) and Rust (Fig. 6b, 6d) benchmarks suggest that CatCoder is able to improve the performance of every tested model.
On the Java benchmark, CatCoder improves the average pass@$1$ score by 13.73%, compared to RepoCoder (Fig. 6a). The improvement is universal, ranging from a significant increase for models like DS-6.7B to a more modest increase for models like CL-13B. Similarly, on the Rust benchmark (Fig. 6b), CatCoder consistently outperforms both In-File and RepoCoder for all LLMs. It improves the average pass@$1$ score by 8.4% compared to RepoCoder. Additionally, the relative performance ranking of the models remains largely consistent across both languages, with CL-13B and DS-6.7B generally at the top, and with CG-7B and ML-8B showing more modest results.
Furthermore, the ablation study results, when generalized across models (Fig. 6c, 6d), are consistent with the findings from RQ2. For both Java and Rust, the -CR variant (without Code Retrieval) consistently leads to the most significant performance degradation, while the -TC variant (without Type Context) shows a smaller but still noticeable drop. On average, across all models on the Java benchmark, the pass@$1$ score decreases by 46.89% without Relevant Code Retrieval, and decreases by 17.17% without Type Context Extraction. This pattern holds for Rust as well. On average, removing Relevant Code Retrieval decreases the pass@1 score by 57.94%, while removing Type Context Extraction leads to an 8.21% decrease.
In conclusion, our results support the generalizability of CatCoder. It consistently enhances the performance of a wide variety of LLMs on both Java and Rust, demonstrating that the architectural benefits of CatCoder are model-agnostic.
@@ 5.3.2. Comparison of different LLMs
The results of our generalizability study also offer insights into the current landscape of LLMs for code:
@@
❶ Although the selected general purpose LLM (i.e., ML-8B) claims to be powerful in many aspects (llama3), it falls behind all selected code-specialized models except for CG-7B in our benchmarks. One example is ML-8B versus CL-7B. Although CL-7B is fine-tuned from an older foundation model (i.e., Llama 2 (touvron2023llama2)), it outperforms the new model ML-8B (i.e., Llama 3 (llama3)). This indicates that code generation tasks benefit a lot from fine-tuning models using high-quality code datasets.
@@
❷ From the trend in Fig. 6, although larger models tend to have a better absolute performance than smaller models, the improvement by CatCoder is larger on smaller models. For instance, the pass@1 score increase for models like ML-8B, CQ-7B, and DS-6.7B is more pronounced than the increase for CL-13B. This suggests that providing high-quality, structured context via frameworks like CatCoder can help smaller or less specialized models bridge the performance gap with larger, more powerful ones.
@@
Answer to RQ3: CatCoder is generalizable to many different LLMs, as it improves the pass@$1$ scores of all selected LLMs, compared with both In-File method and RepoCoder. In addition, the performance of different LLMs on the benchmark provides valuable insights.
5.4. RQ4: Scalability
Figure 7. Detailed latency of the components when constructing prompts for 29 code generation tasks sequentially in the closure-compiler project.
Table 8. Average latency of the components when constructing prompts for 29 code generation tasks sequentially in the closure-compiler project. CR: Relevant Code Retrieval, in Section 3.1; TC: Type Context Extraction, in Section 3.2.
Case CR TC Total
Cold start (Task 0) 38.8s 4.62s 38.8s
Avg. after cold start 1.08s 1.32s 1.56s
Avg. speedup w/ cache 36x 3.5x 25x
In real-world software development, the latency of repository-level code generation tools is crucial, as they inherently process considerable volumes of contextual code in large and complex projects. Significant latencies would negate the advantages of automated assistance and lead to user frustration. Therefore, we propose this research question and investigate the scalability of CatCoder when applied to functions within large-scale code repositories.
5.4.1. Experimental Design
To robustly assess CatCoder’s scalability, the closure-compiler repository [^12] The Closure Compiler is a tool for checking and optimizing JavaScript code. The project is developed by Google and has attracted wide attention. in our benchmark (Section 4.3) is selected for evaluation, by considering the following aspects:
  • This repository is the largest one in our benchmark and comprises approximately 400,000 lines of code, representing a non-trivial, real-world system. Thus, it can present a genuine scalability challenge for CatCoder.
  • It is a well-known and actively developed open-source project. Thus, the evaluation is realistic and can reflect the types of complex dependencies, code structures, and API usage patterns that CatCoder would encounter in practical deployment.
Specifically, our benchmark contains 29 distinct function code generation tasks for the closure-compiler repository. For evaluation, we open this repository, invoke CatCoder on these code generation tasks sequentially, and examine the time consumption of each code generation task. This sequential execution simulates a common developer workflow where a programmer might work on several functions or address multiple issues within the same project during a single coding session.
For time consumption, we focus on the latency of CatCoder’s core components (i.e., Relevant Code Retrieval and Type Context Extraction), while excluding the inference latency of the underlying LLM. This is due to that the time taken for LLM inference is highly dependent on external factors that are largely independent of CatCoder, including the specific LLM chosen, the underlying hardware, and the deployment environment. In addition, the length of prompts constructed by CatCoder does not scale with repository sizes, making the inference latency less important in scalability evaluation.
5.4.2. Analysis of Scalability
Fig. 7 shows the detailed latency of the two components of CatCoder when running each sequential code generation task in the closure-compiler repository. In addition, Table 8 shows the statistics of average latency. It is also worth noting that the two components are independent and can run in parallel. Thus, the total time is bounded by the slower component.
We can notice that CatCoder takes a longer time to process the very first code generation task in a repository that it has not previously analyzed. We refer to the first task as “cold start”. The latency of cold start can be attributed to several setup operations in our approach:
  • For Relevant Code Retrieval (CR), the setup operations mainly involve generating embedding vectors for code chunks, which are crucial for subsequent semantic search.
  • For Type Context Extraction (TC), the operations include the static analyzer’s setup processes, such as parsing the codebase and building representations of type hierarchies and dependencies.
As shown in Table 8, the measured cold start time is 38.8 seconds, with the Relevant Code Retrieval component as the main bottleneck. Considering that the closure-compiler repository is large, and it also takes many IDEs to index a new large repository for a long time, we believe that the cold start overhead of CatCoder is still acceptable for enabling its advanced context-aware code generation capabilities.
Following the initial cold start, CatCoder’s performance for subsequent tasks improves significantly. Table 8 shows the reduction in average latency for tasks after the first one: Relevant Code Retrieval time drops to 1.08s, Type Context Extraction to 1.32s, and the total prompt construction time averages 1.56s. The results indicate an average 36x speedup for Relevant Code Retrieval, 3.5x speedup for Type Context Extraction, and 25x speedup for total time. In this case, after the initial waiting, users do not have to wait for a long time each time to receive a response in an interactive development workflow. This is due to CatCoder’s caching mechanism, which stores intermediate data, including embedding vectors of code chunks and intermediate states of the static analyzer. The cached data is computationally expensive to generate, but generally stable across a sequence of related development tasks within a single coding session. The massive speedup after the cold start demonstrates the effective reuse of cache data.
Answer to RQ4: CatCoder shows practical scalability at a series of code generation tasks within large repositories. In the case of the closure-compiler repository, the caching mechanism of CatCoder is able to introduce a 25x average speedup after the cold start.
@@ 6. Discussion
In this section, we discuss the limitations of our approach and outline potential directions for future work, followed by a discussion of threats to validity.
6.1. Limitations of CatCoder
6.1.1. Limitation in the Underlying LLM
@@
Apart from relevant code and type context, an LLM’s internal capability of understanding general syntax and algorithmic principles is equally essential for correct repository-level code generation. Similar to many prior studies (liu2023repobench; wu2024repoformer; zhang2023repocoder; shrivastava2023repository), CatCoder operates on a frozen LLM. As such, this line of work is inherently limited in enhancing the model’s logical reasoning abilities.
Fig. 8 presents a failure case produced by CatCoder. Although the generated code correctly uses all relevant APIs and closely resembles the ground truth implementation, it fails to pass some unit tests due to missing the logic for handling null pointers (highlighted in green in Fig. 8). This example illustrates that function code generation is inherently complex, and the effective use of repository context does not always guarantee correctness. A possible direction for future work could be fine-tuning LLMs on code repositories to enhance repository-level in-context learning, or adopting reinforcement learning approaches (le2022coderl) to improve reasoning and coding capabilities.
6.1.2. Limitation in Retrieval
In the Relevant Code Retrieval component, we adopt a hybrid retrieval strategy that combines sparse (BM25) and dense (embedding-based) methods to achieve high recall and precision. However, like any information retrieval system, it is not infallible. The retriever may still return incomplete or even misleading context. To some extent, the type context in CatCoder helps mitigate this limitation by supplying accessible API information that the LLM can cross-reference, as illustrated in Section 5.1.2. Possible future work includes enhancing the retrieval and reranking pipeline using powerful LLMs.
6.1.3. Limitation in Static Analysis
Although Java is a statically typed programming language, it supports dynamic features such as reflection. The Java Reflection API enables dynamic class loading, object instantiation, and method invocation based on runtime strings. Despite the capabilities of the static analyzer we employ (i.e., Eclipse JDT.LS, commonly used in IDEs), handling these dynamic features remains challenging. For example, if a target function is related to a field whose class is loaded reflectively at runtime (e.g., via the Class.forName method), the analyzer may fail to resolve the actual class and its members. In such cases, the resulting type context may omit information related to the truly dependent class. Therefore, a limitation is that CatCoder may produce incomplete type context when the repository heavily relies on dynamic features.
@@ 6.2. Threats to Validity
6.2.1. Internal Threats
In general, static analyzers may introduce unsound results, which can, in turn, affect dependent approaches. However, we argue that this is not a concern in CatCoder. Specifically, we use static analysis only to locate types and obtain their definitions and implementations. This form of analysis is lightweight and significantly more accurate than other forms, such as pointer or reflection analysis, as it operates solely on explicit type names in the code. Besides, the tools we use (i.e., Eclipse JDT and rust-analyzer) and the analysis we apply have been integrated into widely used IDEs and code editors (e.g., Eclipse and Visual Studio Code). Therefore, we consider further verification of the analysis results in CatCoder unnecessary, and we trust that our evaluation results remain valid.
6.2.2. External Threats
@@ 7.1. Large Language Models for Code
Large language models for code (i.e., code LLMs) refer to a family of Transformer (vaswani2017attention) models pre-trained or fine-tuned on massive code datasets. Previous work has shown the exceptional performance of code LLMs in various code-related tasks, such as code generation (humaneval), test case generation (tufano2020unit), and program repair (program_repair). In general, pre-trained language models for code can be categorized into the following three types:
@@
    Encoder-only models, such as CodeBERT (feng2020codebert) and GraphCodeBERT (guo2020graphcodebert). These models are pre-trained using Masked Language Modeling (MLM) techniques, and are popular and effective for tasks like classification, where a deep understanding of input data is required.
@@
    Encoder-Decoder models, such as the CodeT5 family (wang2021codet5; wang2023codet5+). These models first encode the input into an internal representation and then decode it into an output sequence.
@@
    Decoder-only models, such as Codex (humaneval), CodeGen (nijkamp2022codegen), StarCoder (li2023starcoder) and CodeLlama (codellama). These models are pre-trained using Next Token Prediction tasks and are powerful in generative tasks.
@@ 7.2. Benchmarks for Repository-Level Code Generation
Iyer et al. (concode) are among the first to highlight the importance of the programmatic context and introduce the task of generating class member functions with natural language descriptions and class contexts. In summary, they construct Concode, a large training dataset consisting of Java classes from online code repositories, and propose an encoder-decoder model for code generation. With the development of large language models, many subsequent repository-level code generation benchmarks have arisen in academia. Du et al. (du2023classeval) make an attempt to evaluate LLMs in the more challenging class-level code generation scenario and manually construct a benchmark named ClassEval, which consists of 100 class-level Python code generation tasks.
Compared with previous benchmarks, for example HumanEval (humaneval) and MBPP (mbpp), ClassEval is more pragmatic. However, it is still limited because, although the methods in the tasks are not standalone, the programmatic contexts are still restricted to very common libraries and methods within the same file. In real-world scenarios, a code repository generally spans a large number of files, making cross-file contexts essential for LLMs to generate the correct code (ding2024crosscodeeval).
Toward more realistic evaluation, subsequent benchmarks such as CoderEval (yu2024codereval), CrossCodeEval (ding2024crosscodeeval), and CoderUJB (zeng2024coderujb) are proposed, all of which are built on a set of real-world open-source repositories and utilize cross-file context for evaluation. A key novelty of CoderEval and CrossCodeEval lies in their ability to support multilingual code generation tasks, while CoderUJB innovatively spans five kinds of practical programming tasks.
As LLMs evolve, Liu et al. (liu2023repobench) point out data leakage and memorization problems, which may affect the integrity and trustworthiness of model evaluation results. To mitigate this issue, they propose RepoBench and regularly update the dataset to keep code samples up-to-date. Similarly, Li et al. (li2024evocodebench) propose EvoCodeBench, which includes an automatic pipeline to evolve itself.
@@ 7.3. Repository-Level Code Generation with LLMs
Many existing approaches for repository-level code generation perform retrieval. Some work, such as (private_library; lu2022reacc), searches for useful information from documentation or external databases. Specifically, Zan et al. (private_library) train a dual-encoder model, which encodes the natural language descriptions and API information, respectively, to retrieve possible APIs from the documentation of the library. Lu et al. (lu2022reacc) propose a retrieval-augmented generation approach named ReAcc, which retrieves similar code from a code database they built. However, these approaches are not flexible and general as their retrieval sources (e.g., API documentation and external databases) may not be available in other repositories.
Other retrieval-based approaches explore retrieving similar code snippets within the repositories. Shrivastava et al. (shrivastava2023repository) propose a framework named Repo-Level Prompt Generator to complete the code for holes in single lines. In the framework, they train a prompt proposal classifier that takes the repository files and a set of prompt proposals and outputs the predicted prompt proposal. Zhang et al. (zhang2023repocoder) propose RepoCoder, which utilizes repository information by iteratively retrieving similar code. In the initial iteration, it uses the natural language descriptions as query input, and in subsequent iterations, it uses previously generated code for queries.
Our work CatCoder combines retrieval-based techniques with a new kind of information called type context, and the comparison with RepoCoder in Section 5 shows that type context serves as essential auxiliary information for statically typed languages. In addition, CatCoder does not require model training or multiple generation iterations, making it more practical in real-world settings.
In addition to the above work, which primarily focuses on prompt engineering, several model training frameworks (wu2024repoformer; deepseekcoder) and decoding strategies (agrawal2024monitor) have been proposed to support repository-level code generation. As our approach CatCoder does not modify the underlying LLM, it can be combined with these frameworks to improve effectiveness.